3. Calculate the wavelength, frequency, and energy of a photon emitted by the following

transitions in a hydrogen atom: (Show your work to receive full credit)

6--->5

A)Wavelength(in nm)

B)Frequency (in s^-1)

C) Energy of photon (in J)
Answer #1

Here photon will be emitted
E = R* (1/nf^2 - 1/ni^2)
R is Rydberg constant. R = 2.18*10^-18 J

E = R* (1/nf^2 - 1/ni^2)
E = 2.18*10^-18* (1/5^2 -   1/6^2)
E = 2.664*10^-20 J

Given:
E = 2.664*10^-20 J

use:
E = h*f
2.664*10^-20 J =(6.626*10^-34 J.s)*f
f = 4.021*10^13 Hz

Given:
E = 2.664*10^-20 J

use:
E = h*c/lambda
2.664*10^-20J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda
lambda = 7.462*10^-6 m
lambda = 7462 nm


Answer:
A) 7462 nm
B) 4.021*10^13 s-1
C) 2.664*10^-20 J